# On Linear Quadratic Potential Games

### v3

@@ 3.3 Decoupled dynamics and information structure
    $Q^{i}_{t+1}=\begin{bmatrix}(Q^{i}_{t+1})_{11}&\dots&(Q^{i}_{t+1})_{1N}\\ \vdots&\ddots&\vdots\\ (Q^{i}_{t+1})_{N1}&\dots&(Q^{i}_{t+1})_{NN}\end{bmatrix}.$
@@
Intuitively, this theorem shows that for two arbitrary agents $i,j$, only the loss parameters corresponding to each other’s state and action should be the same. In contrast, in Proposition 5 part (b), we required that across all agents, the input and state loss parameters be the same (except for those corresponding to $t=0$).
@@
$\frac{\partial J_{x}^{i}(\gamma,x_{0})}{\partial k^{i}_{t}}=$ $\sum_{t^{\prime}=t+1}^{t_{f}}\sum_{h=1}^{N}2(\bar{x}_{t^{\prime}}^{h})^{T}(Q^{i}_{t^{\prime}})_{hi}\frac{\partial\bar{x}_{t^{\prime}}^{i}}{k^{i}_{t}}.$
The second derivatives with respect to $k^{i}_{t}$ and $k^{j}_{\tau}$ are computed by applying the chain rule and product rule such that for $j\in[N]\backslash\{i\}$, and for $\tau\in\{0,\dots,t_{f}-1\}$,
@@
where for $j\in[N]\backslash\{i\}$, and for $\tau\in\{0,\dots,t_{f}-1\}$. Therefore, the second derivative of $J^{i}(\gamma)$ with respect to $k^{i}_{t}$ and $k^{j}_{\tau}$ is as follows
@@
×$\times\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{t_{f}-1}^{i}\big{(}x_{t_{f}-1}^{j}\big{)}^{T}],$
@@
since the derivatives of states and action of agent $j$ are $\frac{\partial\bar{x}^{j}_{t_{f}}}{\partial k_{t_{f}-2}^{j}}=-(A^{j}-b^{j}k^{j}_{t_{f}-1})b^{i}(x^{j}_{t_{f}-2})^{T}$ and $\frac{\partial{u}^{j}_{t_{f}-1}}{\partial k_{t_{f}-2}^{j}}=k_{t_{f}-1}b^{j}(x^{j}_{t_{f}-2})^{T}$. From the pair of time steps $(t_{f}-1,t_{f}-1)$, the condition in (10) is met if and only if
@@
As we discussed in Remark 7, this set of equations may not have a unique solution for $x^{i}$ with a dimension higher than one. Note that the term $(b^{i})^{T}((A^{i})^{t^{\prime}})^{T}(Q_{t_{f}}^{i})_{ij}(A^{j})^{t"}b^{j}$ where $t^{\prime},t"\in\{0,\dots,t_{f}-2\}$ will appear in several equations if other pairs of $(t,\tau)$ are calculated. This notion indicates that conditions on $A^{i}$ and $B^{i}$ might be very restrictive such that the LQ game with decoupled dynamics and decoupled feedback information structure satisfying conditions (C1) and (C2) might be the only game that meets condition (10) for all pairs of $(t,\tau)$.∎ After theoretically deriving non-trivial (non-identical interest) classes of state-feedback potential games, we confirm the existence of practically relevant problems that meet the conditions in Theorem 8.
Example 1 continued: In the formation control games in Example 1, condition (C2) is satisfied since $(R^{i}_{t})_{ij}=0$ for $i\in[N]$ and $j\in[N]\backslash\{i\}$ If the weights are symmetric such that $w^{ij}_{t}=w^{ji}_{t}$, the state loss parameter $(Q^{i}_{t})_{ij}=w^{ij}_{t}$ will be equal to $(Q^{j}_{t})_{ij}=w^{ji}_{t}$; hence, condition (C1) is also met. Thus, from Theorem 8, Example 1 with decoupled feedback information structure is an LQ potential game.
Example 2 continued: The dynamic Cournot game described in Example 2 with the loss function in (6) and the decoupled feedback information structure is also potential. The reason is that this game meets conditions (C1) and (C2) in Theorem 8 as $(R^{i}_{t})_{ji}=p_{t}$, $Q^{i}_{t}=0$ for $t<t_{f}$, and $(Q^{i}_{t_{f}})_{ij}=0$ where $i\in[N]$ and $j\in[N]\backslash\{i\}$.
@@ 4 Properties of decoupled LQ potential games
We first derive the potential function for the game satisfying the conditions of Theorem 8. This potential function corresponds to a single-agent optimal control problem. However, this problem is not a standard LQ control due to the structure of the control policy arising from the decoupled feedback information structure. Given that any optimizer of the potential function will be a Nash equilibrium, we verify the control problem does have an optimal solution. However, we show that characterizing or computing the best response is non-trivial.
@@ 4.1 Existence of a Nash equilibrium
For the state part, in the definition of $Q_{t}$, if $Q_{t}^{hl}$ where $h,l\in[N]\backslash\{i\}$ is replaced by $(Q_{t}^{i})_{hl}$, the matrix $Q^{i}_{t}$ is formed, and consequently, the state part of the $i$th agent loss function at each time step, $(x_{t}-d_{t})^{T}Q_{t}^{i}(x_{t}-d_{t})$, is achieved. Note that $(x^{h}_{t}-d^{h}_{t})^{T}Q_{t}^{hl}(x^{h}_{t}-d^{h}_{t})$ and $(x^{h}_{t}-d^{h}_{t})^{T}(Q^{i}_{t})_{hl}(x^{h}_{t}-d^{h}_{t})$ are independent of the agent $i$’s decision. This means replaced terms were dummy games for agent $i$. For the action part, in the definition of $R_{t}$, if $R_{t}^{hl}$ where $h,l\in[N]\backslash\{i\}$ is replaced by $(R_{t}^{i})_{hl}$, $u_{t}^{T}R_{t}u_{t}$ becomes equal to the action part of $i$th agent loss function at time $t$. Note that $u_{t}^{h}R^{hl}_{t}u_{t}^{l}$ and $u_{t}^{h}(R_{t}^{i})_{hl}u_{t}^{l}$ where $h,l\in[N]\backslash\{i\}$ are independent of agent $i$’s policy and represent dummy games. ∎ In contrast to a classical control problem, the controller must maintain a certain sparsity due to the information structure defined in (iii). Thus, we need to examine the existence of an optimal solution, which corresponds to a Nash equilibrium of the game. In Proposition 10 below, we address the existence of Nash equilibrium and discuss the challenge in its characterization in Proposition 11.
@@
4.2 Challenge in computing the Nash equilbria
Recent works have proven that gradient descent can be applied for computing the optimal control policy in LQ problems [8] and LQ problems with sparsity constraints [9]. For the LQ regulator problem, [8] proved that gradient descent converges to the unique optimum, based on the uniqueness of the global optimum and the gradient-dominated and almost smooth nature of loss function (see Lemmas 3 and 6 in [8]). In [9], this result was extended to finite-horizon structured control problems, a class that encompasses our decoupled LQ potential games. However, the convergence proof in [9] relied on the so-called quadratic invariance property of the sparsity constraints. Unfortunately, as shown below, the quadratic invariance is not generally satisfied for the decoupled information structure considered here.
@@ 4.2 Challenge in computing the Nash equilbria
Consequently, even if we can prove the convergence of the gradient descent to a stationary point, this point is not guaranteed to be a Nash equilibrium. Due to the nonlinear dependence of $k^{i,*}_{t}$ on the policies at previous steps, making assumptions about the uniqueness of the set $(k^{i,*}_{0},\dots,k^{i,*}_{t_{f}-1})$ for which the gradient is zero becomes difficult. The roots of a set of $t_{f}$ non-linear and non-convex equations may represent multiple stationary points, not all of which are the optimal structured control. Thus, while a Nash equilibrium is a stationary point, other stationary points may exist that are not Nash equilibria.
@@ 4.3 Convergence of policy gradient to a stationary point
Note that the term $-\Lambda\nabla\Pi(.)\in\mathbb{R}^{t_{f}n}$ is a descent direction, $-(\nabla\Pi(.))^{T}\Lambda\nabla\Pi(.)<0,$ and $\Pi$ is continuously differentiable and bounded below in every direction. Thus, from Lemma 3.1 in [19], there exists an $\eta$ satisfying the Wolfe conditions, and it can be found using the bisection algorithm presented in [3, Proposition 5.5]. It follows from [19] that $\sum_{m\geq 0}\biggl{(}\frac{\nabla\Pi(k(m))^{T}\Lambda\nabla\Pi(k(m))}{\|\Lambda\nabla\Pi(k(m))\|}\biggr{)}^{2}<\infty.$ This condition is true if and only if $\nabla\Pi(k_{m})$ converges to 00, which implies that: $\lim_{m\to\infty}k(m)=k^{s},$ where $k^{s}$ is a stationary point. Note that variants of policy gradient above, such as stochastic policy gradient assuming no access to the model, can be developed.
@@ 5 Conclusion
Future research can explore approaches to compute the Nash equilibria of the identified potential LQ games, investigate alternative control structures beyond the linear state feedback, and extend the potential LQ game characterization to an infinite-horizon setting.

### v4

@@ 3.3 Decoupled dynamics and information structure
    $Q^{i}_{t+1}=\begin{bmatrix}(Q^{i}_{t+1})_{11}&\dots&(Q^{i}_{t+1})_{1N}\\ \vdots&\ddots&\vdots\\ (Q^{i}_{t+1})_{N1}&\dots&(Q^{i}_{t+1})_{NN}\end{bmatrix}$. (16)
@@
For a clear comparison of the potential conditions in Proposition 5 and Theorem 9, let us write the action part of the loss function in equation (4) in a closed form, $J^{i}_{u}(\gamma,x_{0})=\sum_{t=0}^{t_{f}-1}(u_{t})^{T}R^{i}_{t}u_{t}$ where
$R^{i}_{t}:=\begin{bmatrix}(R^{i}_{t})_{11}&\dots&(R^{i}_{t})_{1N}\\ \vdots&\ddots&\vdots\\ (R^{i}_{t})_{N1}&\dots&(R^{i}_{t})_{NN}\end{bmatrix}$, (17)
and $u=(u^{1},\dots,u^{N})\in\mathbb{R}^{N}$. Note that the state part of the loss function $J^{i}_{x}(\gamma,x_{0})$ in equation (3) can also be written in a closed form when the dynamics are decoupled using the definition of $Q_{t}^{i}$ in equation (16). Proposition 5 states that the game is potential with a full-state feedback information structure if and only if all loss function parameters for both players are the same (except for those corresponding to $t=0$), i.e., $R^{i}_{t}=R^{j}_{t}$ and $Q^{i}_{t}=Q^{j}_{t}$. However, Theorem 9 indicates that the game is potential with a decoupled state feedback information structure if the non-diagonal elements of $R^{i}_{t}$ and $Q^{i}_{t}$ are equal to the non-diagonal ones of $R^{j}_{t}$ and $Q^{j}_{t}$, respectively. It does not impose any conditions on the diagonal elements. As mentioned in Remark 6, the LQ games defined in [2, 25] with non-zero $(R^{i}_{t})_{ii}$ for $i\in[N]$ are assumed to be potential under full-state linear feedback structure if $Q_{t}^{i}=Q_{t}^{j}$ and non-diagonal elements of $R^{i}_{t}$ and $R^{j}_{t}$ are zero. Although the discussion in Remark 6 indicates that this claim is wrong, according to Theorem 9 and assuming decoupled dynamics, one can show that these games are an LQ potential game with decoupled state linear feedback information structure.
@@
$\frac{\partial J_{x}^{i}(\gamma,x_{0})}{\partial k^{i}_{t}}=$ $\sum_{t^{\prime}=t+1}^{t_{f}}\sum_{h=1}^{N}2(\bar{x}_{t^{\prime}}^{h})^{T}(Q^{i}_{t^{\prime}})_{hi}\frac{\partial\bar{x}_{t^{\prime}}^{i}}{\partial k^{i}_{t}}.$
[…]
@@
[…]
@@
$\times\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}[x_{t_{f}-1}^{i}\big(x_{t_{f}-1}^{j}\big)^{T}],$
@@
since the derivatives of states and action of agent $j$ are $\frac{\partial\bar{x}^{j}_{t_{f}}}{\partial k_{t_{f}-2}^{j}}=-(A^{j}-b^{j}k^{j}_{t_{f}-1})b^{j}(x^{j}_{t_{f}-2})^{T}$ and $\frac{\partial{u}^{j}_{t_{f}-1}}{\partial k_{t_{f}-2}^{j}}=k_{t_{f}-1}b^{j}(x^{j}_{t_{f}-2})^{T}$. From the pair of time steps $(t_{f}-1,t_{f}-1)$, the condition in (10) is met if and only if
@@
As discussed in Remark 7, this set of equations may not have a unique solution for $x^{i}$ with a dimension higher than one. Note that the term $(b^{i})^{T}((A^{i})^{t^{\prime}})^{T}(Q_{t_{f}}^{i})_{ij}(A^{j})^{t"}b^{j}$ where $t^{\prime},t"\in\{0,\dots,t_{f}-2\}$ will appear in several equations if other pairs of $(t,\tau)$ are calculated. This notion indicates that conditions on $A^{i}$ and $b^{i}$ might be very restrictive such that the LQ game with decoupled dynamics and decoupled feedback information structure satisfying conditions (C1) and (C2) might be the only game that meets condition (10) for all pairs of $(t,\tau)$.∎
Remark 10.
To investigate the effect of the noisy dynamics on the results of Theorem 9, we add a random noise $w^{i}_{t}$ drawn from the distribution $\mathcal{W}^{i}$ with zero mean to the set of decoupled dynamics in equation (5). By defining $w_{t}:=(w^{1}_{t},\dots,w^{N}_{t})$, the loss function of the LQ game with noisy decoupled dynamics can be written as the one in equation 15. Thus, Lemma 4 is still valid for the game with noisy dynamics and decoupled state linear feedback information structure, and the derivative of $c_{t}(x_{t+1},u_{t})$ with respect to $x_{t}$ or $u_{t}^{j}$ is similar to the game without noise. For decoupled state linear feedback information structure, the derivatives of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $k_{\tau}^{i}$ are zero for $t+1\geq\tau$ and for $t+1<\tau$ can be written as a product of a coefficient that do not depend on $w_{t}$ and $x^{i}_{\tau}$. Note that the derivatives of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to $k_{t}^{j}$ for $j\neq i$ are zero. This observation implies that by investigating conditions in Lemma 4, the potential conditions derived for the LQ game with noisy decoupled dynamics and full-state feedback information structure are the same as those for the game without noise in Theorem 9. Details are provided in Appendix A.
After theoretically deriving non-trivial (non-identical interest) classes of state-feedback potential games, we confirm the existence of practically relevant problems that meet the conditions in Theorem 9.
[…]
@@ 4 Properties of decoupled LQ potential games
We first derive the potential function for the game satisfying the conditions of Theorem 9. This potential function corresponds to a single-agent optimal control problem. However, this problem is not a standard LQ control due to the structure of the control policy arising from the decoupled state linear feedback information structure. Given that any optimizer of the potential function will be a Nash equilibrium, we verify that the control problem has an optimal solution. However, we show that characterizing or computing the best response is non-trivial.
@@ 4.1 Existence of a Nash equilibrium
[…]
@@
4.2 Challenge in computing the Nash equilibria
Recent works have proven that policy gradient can be applied for computing the optimal control policy in LQ problems [8] and LQ problems with sparsity constraints [9]. For the LQ regulator problem, [8] proved that policy gradient converges to the unique optimum, based on the uniqueness of the global optimum and the gradient-dominated and almost smooth nature of the loss function (see Lemmas 3 and 6 in [8]). In [9], this result was extended to finite-horizon structured control problems, a class that encompasses our decoupled LQ potential games. However, the convergence proof in [9] relied on the so-called quadratic invariance property of the sparsity constraints. Unfortunately, as shown below, the quadratic invariance is not generally satisfied for the decoupled information structure considered here.
@@ 4.2 Challenge in computing the Nash equilibria
[…]
@@ 4.3 Convergence of policy gradient to a stationary point
Note that the term $\Lambda\nabla\Pi(.)\in\mathbb{R}^{t_{f}n}$ is a descent direction, $(\nabla\Pi(.))^{T}\Lambda\nabla\Pi(.)<0,$ and $\Pi$ is continuously differentiable and bounded below in every direction. Thus, from Lemma 3.1 in [19], there exists an $\eta$ satisfying the Wolfe conditions, and it can be found using the bisection algorithm presented in [3, Proposition 5.5]. It follows from [19] that $\sum_{m\geq 0}\biggl(\frac{\nabla\Pi(k(m))^{T}\Lambda\nabla\Pi(k(m))}{\|\Lambda\nabla\Pi(k(m))\|}\biggr)^{2}<\infty.$ This condition is true if and only if $\nabla\Pi(k_{m})$ converges to 0, which implies that: $\lim_{m\to\infty}k(m)=k^{s},$ where $k^{s}$ is a stationary point. Note that variants of the policy gradient above, such as stochastic policy gradient assuming no access to the model, can be developed.
4.4 Simulation
In this subsection, we implement the policy gradient method introduced in equation (21). Our goal is to evaluate the convergence of the policy gradient to a stationary point.
We consider a game with $N=4$ agents, where each agent has a state of dimension $n^{i}=1$ and scalar control actions. The time horizon is $t_{f}=10$. To construct the example, we randomly generate the parameters of the game. The values of $A^{i}\in\mathbb{R}$ and $b^{i}\in\mathbb{R}$ are independently sampled from a normal distribution with zero mean and unit variance. Since the matrices $Q_{t}\in\mathbb{R}^{4\times 4}$ and $R_{t}\in\mathbb{R}^{4\times 4}$ of the potential function should be respectively positive semidefinite and positive definite, we first generate the auxiliary matrices $M_{t}\in\mathbb{R}^{4\times 4}$ and $N_{t}\in\mathbb{R}^{4\times 4}$ with entries drawn from the same normal distribution, and then we set $Q_{t}=M_{t}^{\top}M_{t}$ and $R_{t}=N_{t}^{\top}N_{t}+0.01I_{4}$.
Starting from a random $k(0)$, we run the policy iteration for $10^{4}$ steps. We observe that $k(m)$ converges to a stationary point $\tilde{k}$. Then, we repeat the experiment twenty times with different random initializations $k(0)$. The values $k^{i}_{t}(0)$ are randomly sampled from a standard normal distribution. As shown in Fig 1(a), all simulations converge to a stationary point. Furthermore, Fig 1(b) shows that the distance from the previously computed $\tilde{k}$ approaches zero in all simulations, indicating convergence to the same stationary point. As proved in Proposition 12, the game admits at least a Nash equilibrium. Since we found only one stationary point, and any Nash equilibrium has to be a stationary point, it would seem plausible that the $\tilde{k}$ is a Nash equilibrium. However, as discussed in Subsection 4.2, proving that this point is a Nash equilibrium requires demonstrating that no agent can improve their outcome by unilaterally selecting a different policy. Due to the nonlinear dependence of the optimal strategy $k^{i,*}_{t}$ on the policies at both previous and successive steps, checking for potential improvements would necessitate an exhaustive search over the entire policy space. Specifically, this involves exploring a space with dimensions $n\times t_{f}=40$ in our current setting.
(a) $|\nabla\Pi(k(m))|$
(b) |k(m)k~||k(m)-$\tilde${k}|
Figure 1: Convergence of policy gradient with the update rule in equation (21) to a stationary point $\tilde{k}$ for 20 simulations. In these figures, $m$ indicates the iteration number, and the abscissa and the ordinate are on a logarithmic scale.
@@ 5 Conclusion
Future research can explore approaches to compute the Nash equilibria of the identified potential LQ games, investigate alternative control structures beyond the decoupled state linear feedback, and extend the potential LQ game characterization to an infinite-horizon setting.
@@
Appendix A Additive Noise
Here, we prove that the results in Proposition 5 and Theorem 9 can be generalized to the LQ game with noise. For this purpose, we introduce the loss function of the game with noisy dynamics, compare the derivatives in condition (10) of Lemma 4 for this loss function and the one without noise, and illustrate that the final potential conditions in Proposition 5 and Theorem 9 are the same for both loss functions. We assume the dynamics of the game include noise as follows.
$x_{t+1}=Ax_{t}+\sum_{i=1}^{N}B^{i}u^{i}_{t}+w_{t},\quad x_{0}\sim\mathcal{D},\ w_{t}\sim\mathcal{W}$ (22)
where $w_{t}$ is a random noise that is assumed to be drawn from the distribution $\mathcal{W}$ with zero mean. The loss function will be as follows.
$J^{i}(\gamma)=$ $\mathop{\mathbb{E}}_{x_{0}\sim\mathcal{D}}\bigl[\sum_{t=0}^{t_{f}-1}\mathop{\mathbb{E}}_{w_{t}\sim\mathcal{W}}c_{t}(x_{t+1},u_{t})\bigr],$ (23)
where the stage loss $c_{t}(x_{t+1},u_{t})$ is defined as follows.
$c_{t}(x_{t+1},u_{t}):=$ $(x_{t+1}-d_{t+1})^{T}Q_{t+1}^{i}(x_{t+1}-d_{t+1})$
$+\sum_{j=1}^{N}\sum_{h=1}^{N}(u_{t}^{j})^{T}(R^{i}_{t})_{jh}u^{h}_{t},$ (24)
Note that for decoupled dynamics, we add a random noise $w^{i}_{t}$ drawn from the distribution $\mathcal{W}^{i}$ with zero mean to the set of decoupled dynamics in equation (5). By defining $w_{t}:=(w^{1}_{t},\dots,w^{N}_{t})$, the loss function of the LQ game with noisy decoupled dynamics can be written as the one in equation 23. Considering the loss function defined above, the definitions of Nash equilibrium, potential game, identical interest games for LQ games, (Definitions 12, and 3) the game pseudo-gradient, $\mathcal{G}(\gamma)$, and its Jacobian in equation (9) do not change. Hence, Lemma 4 is still valid for the game with noisy dynamics. Note that the derivative of $c_{t}(x_{t+1},u_{t})$ with respect to $x_{t}$ or $u_{t}^{j}$ is similar to the game without noise. Therefore, we only investigate the derivative of $x_{t+1}$ and $u_{t}^{i}$ with respect to the decision variable $\gamma_{\tau}^{j}$ in the presence of the noise.
Proposition 5 part (a): For the open-loop information structure, note that the decision-making variable is $\gamma_{t}^{j}=u_{t}^{j}$, and the game dynamics can be written as $x_{t+1}=a^{t}x_{0}+\sum_{\tau=0}^{t}a^{t-\tau}[b^{1}u^{1}_{\tau}+b^{2}u^{2}_{\tau}+w_{\tau}]$. Thus, the derivative of $x_{t+1}$ and $u_{t}^{i}$ with respect to $u_{\tau}^{j}$ is as follows.
$\frac{\partial x_{t+1}}{\partial u_{\tau}^{j}}=$ $\begin{cases}A^{t-\tau}B^{i}&t\geq\tau\\ 0&t<\tau\end{cases}$
$\frac{\partial u_{t}^{i}}{\partial u_{\tau}^{j}}=$ $\begin{cases}1&t=\tau\ \text{and}\ i=j\\ 0&\text{otherwise}\end{cases}$
The derivatives above are similar to the ones without noise. This proves that the conditions in Proposition 5, part (a), are valid for the game with noise.
Proposition 5 part (b): For the full-state feedback information structure, note that the decision-making variable is $\gamma_{t}^{j}=K_{t}^{j}$. The derivative of $x_{t+1}$ and $u_{t}^{i}$ with respect to $K_{\tau}^{j}$ can be derived using the chain rule and assuming the states are scalar (as mentioned in Example 3) as follows.
$\frac{\partial x_{t+1}}{\partial K_{\tau}^{j}}=$ $\begin{cases}-\prod_{t^{\prime}=\tau+1}^{t}\frac{\partial x_{t^{\prime}+1}}{\partial x_{t^{\prime}}}b^{j}x_{\tau}&t>\tau\\ -b^{j}x_{t}&t=\tau\\ 0&t<\tau\end{cases}$
$\frac{\partial u_{t}^{i}}{\partial K_{\tau}^{j}}=$ $\begin{cases}-K_{t}^{i}\frac{\partial x_{t}}{\partial K_{\tau}^{j}}&t>\tau\\ -x_{t}&t=\tau\ \text{and}\ i=j\\ 0&\text{otherwise}\end{cases}$
Note that $\frac{\partial x_{t+1}}{\partial x_{t}}$ is not a function of $w_{t}$; therefore, the derivatives above are similar to the ones without noise. This proves that the conditions in Proposition 5, part (b), are valid for the game with noise.
Theorem 9: For the decoupled state feedback information structure, note that the decision-making variable is $\gamma_{t}^{j}=k_{t}^{j}$. The derivative of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to $k_{\tau}^{i}$ can be derived using the chain rule as follows.
$\frac{\partial x^{i}_{t+1}}{\partial k_{\tau}^{i}}=$ $\begin{cases}-\prod_{t^{\prime}=\tau+1}^{t}\frac{\partial x^{i}_{t^{\prime}+1}}{\partial x^{i}_{t^{\prime}}}b^{i}(x^{i}_{\tau})^{T}&t>\tau\\ -b^{i}(x^{i}_{t})^{T}&t=\tau\\ 0&t<\tau\end{cases}$
$\frac{\partial u_{t}^{i}}{\partial k_{\tau}^{i}}=$ $\begin{cases}-k_{t}^{i}\frac{\partial x^{i}_{t}}{\partial k_{\tau}^{i}}&t>\tau\\ -(x^{i}_{t})^{T}&t=\tau\\ 0&t<\tau\end{cases}$
Note that the derivative of $x^{i}_{t+1}$ and $u_{t}^{i}$ with respect to $k_{t}^{j}$ for $j\neq i$ are zero, and $\frac{\partial x^{i}_{t+1}}{\partial x^{i}_{t}}$ is not a function of $w_{t}$. Hence, the derivatives above are similar to the ones without noise. These results indicate that the conditions derived for the game with the decoupled dynamic defined in equation (5) and a decoupled state feedback information structure in (iii) to be potential do not depend on noise, and consequently, adding the noise does not affect the results.